Grade 11Electric Current

If Ec= 20sin(10^10)pi*t and Em= 10sin400pi*t
are the carrier signal and modulation signal respectively, then what will be modulation index
[a]50% [b]60%
[c]64% [d]48%

Aniket debnath
7 Years agoGrade 11

1 Answer

Khimraj
7 Years ago
modulation index = (Vm/Vc)*100 = (10/20)*100 = 50%
Hope it clears.